# EgoBrain: Synergizing Minds and Eyes For Human Action Understanding

### v1

@@ 1 Introduction
The explosive growth of artificial intelligence has greatly advanced the field of Brain-computer interfaces (BCI), with massive research efforts to understand brain functions from neural recordings. Among various neural signals, non-invasive systems such as scalp electroencephalograph (EEG) are more scalable, cost-effective, and safer for large-scale adoption willett:nature2021 ; anumanchipalli:nature2019 ; sivasakthivel:sr2025 ; metzger:nature2023 ; bai2023dreamdiffusion ; li2025translating ; lan2023seeing , thus appealing increasing interest to connect EEG with human perceptions and intentions. Boosted by deep learning techniques, booming breakthroughs have been seen in recent years to decode visual and acoustic stimuli in controlled laboratory settings. For example, recent works achieved accuracies of 15.6% in a 200-way zero-shot task on the EEG-image dataset song2023decoding and 21.9% in a 9-way task on the EEG-video dataset liu:nips2024 . However, the visual stimuli in existing studies were merely presented on screens and the informative environmental background was ignored. Moreover, the active interactions between the subjects and the environment are less explored due to the passive settings in the experiments.
To better capture real-world human perceptions and actions, we introduce egocentric (first-person-view) vision as a complementary modality to EEG. The egocentric vision has emerged as a powerful paradigm for modeling human-object interactions and perceptual processes in real-world settings, with representative large-scale datasets such as EPIC-KITCHENSdamen:ijcv21 , Ego4Dgrauman:cvpr22 and HoloAssistwang:iccv23 . These datasets primarily capture observable outcomes from a human-like perspective, yielding valuable analysis of human behavior such as action recognition, hand pose estimation and human-object interaction understanding.
@@
Similar to other multimodal tasks with synchronized timeline, it’s crucial to handle the shared temporal structure carefully and fuse information from modalities for downstream prediction. Upon our EgoBrain dataset, we present an adaptive Brain-Time Interval Machine (Brain-TIM) model, inspired from chalk:cvpr2024 to integrate synchronized visual and EEG signals and capture rich multimodal information for action understanding. Each modality is processed through modality-specific embedding layers and merged to the aggregated global context, while the shared temporal structure is explicitly modeled using the Time Interval MLP (TIM) module. We then conducted experiments with our Brain-TIM to evaluate both the standalone effectiveness of individual modalities and their synergy, and the highlighted results confirmed that the fusion of EEG and vision consistently outperforms unimodal approaches across multiple experiments. Further visualization provide deeper insight into the complementary roles of egocentric vision and EEG signals. We summarize our contributions as:
@@ 2 Related Work
EEG & Vision Integration: In recent years, combining electroencephalography (EEG) with visual data has emerged as a central theme in brain–computer interface (BCI) research, elucidating cognitive processes and motor intentionsmushtaq:nhb2024 ; guttmann:scidata2025 ; bertoni:nc2025 ; dreyer:scidata2023 ; kaya:scidata2018 . EEG’s high temporal resolution and portability enable real-time monitoring of brain states, yet most work examines resting-state responses to static visual stimuli, neglecting neural dynamics during natural movementyang:scidata2025 ; liu:scidata2025 ; ma:scidata2022 ; ma:scidata2020 ; liu:scidata2024 . A few studies have recorded EEG during active locomotion—for example, assessing cognitive load while walking in a lower-limb exoskeletonortiz:scidata2023 —and virtual-reality tasks like supernumerary thumb control via motor imageryalsuradi:scireport2024 . However, these efforts target prosthetic control and lack a systematic exploration of real-world, first-person multimodal interactions in unconstrained movement.
Egocentric Vision Datasets: Recent egocentric video corpora have advanced human–object interaction modeling through varied contexts and annotationsdamen:ijcv21 ; grauman:cvpr22 ; wang:iccv23 ; darkhalil:nips22 ; kwon:iccv21 ; liu:cvpr2022 ; ragusa:wacv21 ; sener:cvpr22 ; ohkawa:cvpr23 ; zhang:eccv2022 ; grauman:cvpr24 ; huang:cvpr2024 . EPIC-KITCHENSdamen:ijcv21 offers detailed kitchen-activity labels, while Ego4Dgrauman:cvpr22 provides the largest in-the-wild egocentric set for 3D perception and social analysis. HoloAssistwang:iccv23 enables multi-user task completion, and Assembly-seriessener:cvpr22 ; ohkawa:cvpr23 and H2Okwon:iccv21 cover procedural and two-hand manipulations. More recent datasets like EgoExoLearnhuang:cvpr2024 and Ego-Exo4Dgrauman:cvpr24 deliver asynchronous and dual-perspective recordings of skilled activities. Unlike these vision-only resources, our EgoBrain jointly captures first-person video and EEG, facilitating cross-modal action recognition.
@@ 4.1 Task Definition
We consider a time-synchronized pair of raw data: the egocentric video stream and the EEG signal sequence, both sharing a common timeline $\mathcal{T}=[0,T]$. The video stream is represented as $V^{\text{raw}}=\{v_{t}\in\mathbb{R}^{H\times W\times 3}\}_{t=0}^{T\cdot f^{v}}$, sampled at a frame rate $f^{v}$, and the EEG signal as $B^{\text{raw}}=\{b_{t}\in\mathbb{R}^{C}\}_{t=0}^{T\cdot f^{b}}$, recorded at $f^{b}$ Hz, where $C$ is the number of channels. The target of action recognition task can be formulated as finding the best mapping from input to the action and verb categories $\hat{y}=f_{\theta}(V,B)\in\{1,\dots,N_{c}\}^{Q}$, where $N_{c}$ equals to 10 for verb classification or 29 for action categories, and $Q$ is the number of consecutive queries which evenly divides the whole time interval $[0,T]$, i.e, the $i$-th query corresponds to the action within time $[(i-1)T/Q,iT/Q]$.
@@ 4.2 Overview of Brain-TIM
An overview of Brain-TIM is presented in Fig. 3. We first extract feature representations for each modality using pre-trained backbone networks tong:nips2022 ; jiang:iclr2024 into $\phi^{v}$ and $\phi^{b}$ (Appendix  A.1). These features are then projected into a shared embedding space via modality-specific embedding layers: $g^{v}$ and $g^{b}$ (Appendix  A.2). The embeddings from different modalities are concatenated to form a unified input sequence for the Transformer module on the right side of the Fig. 3 (Appendix  A.2). Eventually, the Transformer encoder models the temporal dependencies and cross-modal interactions within the sequence (Appendix A.3) and a linear classifier maps the encoded features to the final action category predictions (Appendix A.4).
@@ 4.3 Sequence Concatenation with Temporal and Modality-Specific Information
To discriminate tokens before feeding them into transformer and enhance the representations with temporal and modality-specific information, we introduce the Time-aware Token Embedding and the Modality-specific embedding following chalk:cvpr2024 .
@@
For the feature tokens $\tilde{\mathbf{e}}_{i}^{v}$ and $\tilde{\mathbf{e}}_{i}^{b}$, the time interval is determined by the $i$-th window $[t_{i},t_{i}+\Delta t)$, also named feature time in chalk:cvpr2024 . The temporal embedding is thus calculated as $\mathbf{e}_{i}^{f}=I(t_{i},t_{i}+\Delta t)\in\mathbb{R}^{D}$. As for the CLS tokens $\mathbf{c}^{v}_{j}$ and $\mathbf{c}^{b}_{j}$ , the time interval corresponds to the $j$-th query $[(j-1)T/Q,jT/Q]$, also known as query time. And the temporal embedding is similarly obtained as $\mathbf{e}_{j}^{q}=I((j-1)T/Q,jT/Q)\in\mathbb{R}^{D}$.
@@
$\mathbf{X}=\text{Concat}\big{(}\underbrace{\{\tilde{\mathbf{e}}^{v}_{i}\|\mathbf{e}^{f}_{i}+\mathbf{m}^{v}\}_{i=1}^{N}}_{\text{visual feature block}},\ \underbrace{\{\tilde{\mathbf{e}}^{b}_{i}\|\mathbf{e}^{f}_{i}+\mathbf{m}^{b}\}_{i=1}^{N}}_{\text{brain feature block}},\ \underbrace{\{\mathbf{c}_{j}^{v}\|\mathbf{e}^{q}_{j}+\mathbf{m}^{v}\}_{j=1}^{Q}}_{\text{visual CLS token block}},\ \underbrace{\{\mathbf{c}_{j}^{b}\|\mathbf{e}^{q}_{j}+\mathbf{m}^{b}\}_{j=1}^{Q}}_{\text{brain CLS token block}}\big{)},$
@@ 5.1 Action Classification Results on EgoBrain
Table 1: Action recognition results on validation set. We present the baseline performance on action classification of our EgoBrain via Brain-TIM. The validation set follows a cross-subject, within-scene protocol and does not involve scene variation compared to the training set. The primary evaluation metric (Action Acc. % $\rightarrow$ @Top-1) is reported with the best-performing results highlighted in bold.
Modality Fusion Cross-subject Only Cross-subject & Cross-scene
Verb Acc.Action Acc.Verb Acc.Action Acc.%
Brain Only - $23.19\pm 0.29$ $12.12\pm 0.43$ $19.41\pm 1.57$ $9.36\pm 0.52$
Visual Only - $92.50\pm 0.73$ $82.80\pm 0.48$ $81.67\pm 1.89$ $63.40\pm 0.95$
Visual & Brain Spatial $92.80\pm 0.42$ $82.91\pm 0.31$ $83.74\pm 0.62$ $64.81\pm 1.04$
Visual & Brain Brain-TIM $92.72\pm 0.52$ 84.48 $\pm$ 0.23 $83.43\pm 0.41$ 66.70 $\pm$ 0.83
Unimodal Comparison: As a well-studied computer vision problem, the visual modality demonstrates significantly strong performance, achieving 92.50% Top-1 accuracy for action classification and 82.80% for verb classification in the cross-subject setting. Thanks to its superior spatial resolution and contextual richness, egocentric visual input provides fine-grained cues that are critical for distinguishing actions and achieved the performances of 83.43% and 66.70% for verb and action classification even under the challenging cross-subject and cross-scene settings.
@@
Multmodal Comparison: As shown in Tab. 1, fusing EEG with visual inputs boosts most of the accuracies by a significant margin. Specifically, on the most difficult cross-scene setting, Brain-TIM with both modalities as inputs yields a 3.30% absolute improvement over the vision-only baseline on the 29-class action recognition task. This performance gain answers RQ1 and highlights the semantic complementarity between the two modalities: while the visual stream captures external manifestations of action, EEG encodes neural signatures of motor intention and implicit knowledge that are not observable from eyes alone.
@@
Action Acc.@Top-1 Mean Action Acc.@Top-1 STD 7.44 0.39 7.54 0.05 6.15 0.06 9.36 0.52 64.94 3.64 64.67 1.75 64.01 6.34 63.40 0.95 65.71 0.43 65.81 2.15 66.39 0.09 66.18 2.22 66.70 0.83
5.2 Ablation Study
Temporal fusion vs. Spatial fusion: Our Brain-TIM generally uses the temporal fusion for the two modalities to construct the input X. A simpler approach could be directly concatenating along the feature dimension, also known as the spatial fusion. In this way, the input $X$ is constructed as:
$\mathbf{X}=\text{Concat}\big{(}\underbrace{\{\tilde{\mathbf{e}}^{v}_{i}\|\tilde{\mathbf{e}}^{b}_{i}\|\mathbf{e}^{f}_{i}\}_{i=1}^{N}}_{\text{ feature block}},\ \underbrace{\{\mathbf{c}_{j}^{v}\|\mathbf{c}_{j}^{b}\|\mathbf{e}^{q}_{j}\}_{j=1}^{Q}}_{\text{CLS token block}}\big{)}\in\mathbb{R}^{(N+Q)\times 3D}.$
We examined the arbitration between fusion choice and also present the result in Tab. 1. Although the two fusion methods are comparable on the verb classification, our Brain-TIM with temporal fusion significantly outperforms the spatial fusion for the more complex action recognition. This result indicates that while the simple feature concatenation works in a rough classification scenario, modeling temporal dependencies across time steps leads to more robust and discriminative multimodal representations, particularly improving performance on complex real-world action recognition tasks.
Effectiveness of components: Table 2 presents the performance under three modality settings: Brain Only, Visual Only, and Visual & Brain. For each setting, we evaluate different combinations of three key components: the embedding layer $g^{(v,b)}$, the time interval MLP $I^{(v,b)}$, and the modality embedding $\mathbf{m}^{(v,b)}$.
@@ 5.3 Detailed Analysis
Benefits of Integrating EEG: As shown in Fig.5a, when a participant is drawing in a notebook, the visual model misclassifies the action as “writing” due to the high visual similarity between the two tasks. However, after incorporating EEG data, the model correctly classifies the action as “drawing”. This suggesting that EEG signals may capture neural patterns related to task intent and offer additional discriminative cues. This suggests that EEG captures distinct neural activations associated with visuospatial motor planning as opposed to language-related tasks—a distinction that has been well documented in prior neuroscience studiestang:tim2024 .
Another example in Fig.5a involves the action of “drink - bitter ground juice”. Due to occlusion, the subject’s hand and the cup are not visible, and the visual model misclassifies the action as “snack” based on nearby contextual visual cues (e.g., a bag of chips). With EEG integration, the model not only correctly identifies the verb “drink” but also the object “bitter melon juice”. This improvement likely stems from EEG’s ability to reflect orofacial motor patterns and anticipatory neural activity related to swallowing, which differ from those associated with chewing saito:dysphagia2024 . The results underscore EEG’s value in disambiguating semantically similar actions when vision is limited or misleading.
Limitations of Integrating EEG: Despite these advantages, EEG does not always lead to improved recognition. Fig.5b presents failure cases of the multimodal model. The participant is actually operating PowerPoint, but the model incorrectly identifies the task as using a drawing application. Possible reason is that the participant is creating multiple rectangles, evoking visuomotor activity patterns that resemble those during freehand drawing. Prior studies have shown that such overlapping cognitive strategies lead to similar EEG signatures dvorak:tnsre2018 , making semantic discrimination harder.

### v2

@@ 1 Introduction
The explosive growth of artificial intelligence has greatly advanced the field of Brain-computer interfaces (BCI), with massive research efforts to understand brain functions from neural recordings. Among various neural signals, non-invasive systems such as scalp electroencephalograph (EEG) are more scalable, cost-effective, and safer for large-scale adoption (Willett et al., 2021; Anumanchipalli et al., 2019; Sivasakthivel et al., 2025; Metzger et al., 2023; Bai et al., 2023; Li et al., 2025; Lan et al., 2023), thus appealing increasing interest to connect EEG with human perceptions and intentions. Boosted by deep learning techniques, booming breakthroughs have been seen in recent years to decode visual and acoustic stimuli in controlled laboratory settings. For example, recent works achieved accuracies of 15.6% in a 200-way zero-shot task on the EEG-image dataset (Song et al., 2023) and 21.9% in a 9-way task on the EEG-video dataset (Liu et al., 2024b). However, the visual stimuli in existing studies were merely presented on screens and the informative environmental background was ignored. Moreover, the active interactions between the subjects and the environment are less explored due to the passive settings in the experiments.
To better capture real-world human perceptions and actions, we introduce egocentric (first-person-view) vision as a complementary modality to EEG. The egocentric vision has emerged as a powerful paradigm for modeling human-object interactions and perceptual processes in real-world settings, with representative large-scale datasets such as EPIC-KITCHENS(Damen et al., 2022), Ego4D(Grauman et al., 2022) and HoloAssist(Wang et al., 2023). These datasets primarily capture observable outcomes from a human-like perspective, yielding valuable analysis of human behavior such as action recognition, hand pose estimation and human-object interaction understanding.
@@
Similar to other multimodal tasks with synchronized timeline, it’s crucial to handle the shared temporal structure carefully and fuse information from modalities for downstream prediction. Upon our EgoBrain dataset, we present an adaptive Brain-Time Interval Machine (Brain-TIM) model, inspired from (Chalk et al., 2024) to integrate synchronized visual and EEG signals and capture rich multimodal information for action understanding. Each modality is processed through modality-specific embedding layers and merged to the aggregated global context, while the shared temporal structure is explicitly modeled using the Time Interval MLP (TIM) module. We then conducted experiments with our Brain-TIM to evaluate both the standalone effectiveness of individual modalities and their synergy, and the highlighted results confirmed that the fusion of EEG and vision consistently outperforms unimodal approaches across multiple experiments. Further visualization provide deeper insight into the complementary roles of egocentric vision and EEG signals. We summarize our contributions as:
@@ 2 Related Work
EEG & Vision Integration: In recent years, combining electroencephalography (EEG) with visual data has emerged as a central theme in brain–computer interface (BCI) research, elucidating cognitive processes and motor intentions(Mushtaq et al., 2024; Guttmann-Flury et al., 2025; Bertoni et al., 2025; Dreyer et al., 2023; Kaya et al., 2018). EEG’s high temporal resolution and portability enable real‐time monitoring of brain states, yet most work examines resting‐state responses to static visual stimuli, neglecting neural dynamics during natural movement(Yang et al., 2025; Liu et al., 2025; Ma et al., 2022; 2020; Liu et al., 2024a). A few studies have recorded EEG during active locomotion—for example, assessing cognitive load while walking in a lower‐limb exoskeleton(Ortiz et al., 2023)—and virtual‐reality tasks like supernumerary thumb control via motor imagery(Alsuradi et al., 2024). However, these efforts target prosthetic control and lack a systematic exploration of real‐world, first‐person multimodal interactions in unconstrained movement.
Egocentric Vision Datasets: Recent egocentric video corpora have advanced human–object interaction modeling through varied contexts and annotations(Damen et al., 2022; Grauman et al., 2022; Wang et al., 2023; Darkhalil et al., 2022; Kwon et al., 2021; Liu et al., 2022; Ragusa et al., 2021; Sener et al., 2022; Ohkawa et al., 2023; Zhang et al., 2022; Grauman et al., 2024; Huang et al., 2024). EPIC‐KITCHENS(Damen et al., 2022) offers detailed kitchen‐activity labels, while Ego4D(Grauman et al., 2022) provides the largest in‐the‐wild egocentric set for 3D perception and social analysis. HoloAssist(Wang et al., 2023) enables multi‐user task completion, and Assembly‐series(Sener et al., 2022; Ohkawa et al., 2023) and H2O(Kwon et al., 2021) cover procedural and two‐hand manipulations. More recent datasets like EgoExoLearn(Huang et al., 2024) and Ego‐Exo4D(Grauman et al., 2024) deliver asynchronous and dual‐perspective recordings of skilled activities. Unlike these vision‐only resources, our EgoBrain jointly captures first‐person video and EEG, facilitating cross‐modal action recognition.
@@ 4.1 Task Definition
We consider a time-synchronized pair of raw data: the egocentric video stream and the EEG signal sequence, both sharing a common timeline $\mathcal{T}=[0,T]$. The video stream is represented as $V^{\text{raw}}=\{v_{t}\in\mathbb{R}^{H\times W\times 3}\}_{t=0}^{T\cdot f^{v}}$, sampled at a frame rate $f^{v}$, and the EEG signal as $B^{\text{raw}}=\{b_{t}\in\mathbb{R}^{C}\}_{t=0}^{T\cdot f^{b}}$, recorded at $f^{b}$ Hz, where $C$ is the number of channels. The target of action recognition task can be formulated as finding the best mapping from input to the action and verb categories $\hat{y}=f_{\theta}(V,B)\in\{1,\dots,N_{c}\}^{Q}$, where $N_{c}$ equals to 10 for verb classification or 29 for action categories, and $Q$ is the number of consecutive queries which evenly divides the whole time interval $[0,T]$, i.e, the $i$-th query corresponds to the action within time [(i1)T/Q,iT/Q][(i-1)T/Q,iT/Q].
@@ 4.2 Overview of Brain-TIM
An overview of Brain-TIM is presented in Fig. 3. We first extract feature representations for each modality using pre-trained backbone networks (Tong et al., 2022; Jiang et al., 2024) into $\phi^{v}$ and $\phi^{b}$ (Appendix  A.1). These features are then projected into a shared embedding space via modality-specific embedding layers: $g^{v}$ and $g^{b}$ (Appendix  A.2). The embeddings from different modalities are concatenated to form a unified input sequence for the Transformer module on the right side of the Fig. 3 (Appendix  A.2). Eventually, the Transformer encoder models the temporal dependencies and cross-modal interactions within the sequence (Appendix A.3) and a linear classifier maps the encoded features to the final action category predictions (Appendix A.4).
@@ 4.3 Sequence Concatenation with Temporal and Modality-Specific Information
To discriminate tokens before feeding them into transformer and enhance the representations with temporal and modality-specific information, we introduce the Time-aware Token Embedding and the Modality-specific embedding following (Chalk et al., 2024).
@@
For the feature tokens $\tilde{\mathbf{e}}_{i}^{v}$ and $\tilde{\mathbf{e}}_{i}^{b}$, the time interval is determined by the $i$-th window $t_{i},t_{i}+\Delta t)$, also named feature time in (Chalk et al., 2024). The temporal embedding is thus calculated as $\mathbf{e}_{i}^{f}=I(t_{i},t_{i}+\Delta t)\in\mathbb{R}^{D}$. As for the CLS tokens $\mathbf{c}^{v}_{j}$ and $\mathbf{c}^{b}_{j}$ , the time interval corresponds to the $j$-th query [(j1)T/Q,jT/Q][(j-1)T/Q,jT/Q], also known as query time. And the temporal embedding is similarly obtained as $\mathbf{e}_{j}^{q}=I((j-1)T/Q,jT/Q)\in\mathbb{R}^{D}$.
@@
[…]
@@ 5.1 Action Classification Results on EgoBrain
Table 1: Action recognition results on the EgoBrain test set. We systematically evaluate unimodal (Brain only, Visual only) and multimodal (Visual+Brain) models under two protocols: cross-subject only and cross-subject & cross-scene. The table reports the parameter scale (Params) of each model and the mean $\pm$ standard deviation across five random seeds to ensure statistical reliability. The primary evaluation metric is Top-1 accuracy (%), with the best results highlighted in bold.
Protocol Modality Encoder Params Verb Acc.% Action Acc.%
Cross-subject
only
Brain only
LaBraM
(Jiang et al., 2024)
5.8M $21.53\pm 0.99$ $8.44\pm 2.25$
Visual only
VideoMAE
(Tong et al., 2022)
305.0M $88.95\pm 0.80$ $78.44\pm 0.71$
Visual + Brain
VideoMAE + LaBraM
(Tong et al., 2022; Jiang et al., 2024)
310.8M 90.11 $\pm$ 1.10 80.16 $\pm$ 1.67
Cross-subject &
Cross-scene
Brain only
LaBraM
(Jiang et al., 2024)
5.8M $19.41\pm 1.57$ $9.36\pm 0.52$
Visual only
VideoMAE
(Tong et al., 2022)
305.0M $81.67\pm 1.89$ $63.40\pm 0.95$
Visual + Brain
VideoMAE + LaBraM
(Tong et al., 2022; Jiang et al., 2024)
310.8M 83.43 $\pm$ 0.41 66.70 $\pm$ 0.83
Unimodal Comparison: As a well-studied computer vision problem, the visual modality demonstrates significantly strong performance, achieving 88.95% Top-1 accuracy for verb classification and 78.44 % for action classification in the cross-subject setting. Thanks to its superior spatial resolution and contextual richness, egocentric visual input provides fine-grained cues that are critical for distinguishing actions and achieved the performances of 81.67% and 63.40% for verb and action classification even under the challenging cross-subject and cross-scene settings.
@@
Multmodal Comparison: As shown in Tab. 1, fusing EEG with visual inputs boosts accuracies across both evaluation protocols. Under the cross-subject only setting, the vision-only baseline achieves 78.44% Top-1 accuracy for action classification, while Brain-TIM reaches 80.16%, giving a 1.72% improvement. Specifically, under the most difficult cross-scene setting, the vision-only baseline achieves 63.40% Top-1 accuracy for action classification, while Brain-TIM with both modalities reaches 66.70%, yielding a 3.30% absolute improvement on the 29-class task. This performance gain answers RQ1 and further highlights the semantic complementarity between the two modalities: while the visual stream captures external manifestations of action, EEG encodes neural signatures of motor intention and implicit knowledge that are not observable from eyes alone.
Importantly, the performance gains of EEG are not attributable to differences in model capacity, but rather to its unique cognitive value. Our EEG encoder is extremely lightweight, with only 5.8M parameters (approximately 1/52 of the visual backbone), yet it still delivers statistically significant improvements. This demonstrates that EEG provides indispensable complementary information in cases of visual ambiguity or occlusion, thereby rendering visual understanding more complete.
5.2 Ablation Study
Table 2 presents the performance under three modality settings: Brain Only, Visual Only, and Visual & Brain. For each setting, we evaluate different combinations of three key components: the embedding layer $g^{(v,b)}$, the time interval MLP $I^{(v,b)}$, and the modality embedding $\mathbf{m}^{(v,b)}$.
@@ 5.2 Ablation Study
Action Acc.% @Top-1 Mean
Action Acc.% @Top-1 STD
7.44
0.39
7.54
0.05
6.15
0.06
9.36
0.52
64.94
3.64
64.67
1.75
64.01
6.34
63.40
0.95
65.71
0.43
65.81
2.15
66.39
0.09
66.18
2.22
66.70
0.83
@@ 5.3 Detailed Analysis
Benefits of Integrating EEG: As shown in Fig.5a, when a participant is drawing in a notebook, the visual model misclassifies the action as “writing” due to the high visual similarity between the two tasks. However, after incorporating EEG data, the model correctly classifies the action as “drawing”. This suggesting that EEG signals may capture neural patterns related to task intent and offer additional discriminative cues. This suggests that EEG captures distinct neural activations associated with visuospatial motor planning as opposed to language-related tasks—a distinction that has been well documented in prior neuroscience studies(Tang et al., 2024).
@@
Another example in Fig.5a involves the action of “drink - bitter ground juice”. Due to occlusion, the subject’s hand and the cup are not visible, and the visual model misclassifies the action as “snack” based on nearby contextual visual cues (e.g., a bag of chips). With EEG integration, the model not only correctly identifies the verb “drink” but also the object “bitter melon juice”. This improvement likely stems from EEG’s ability to reflect orofacial motor patterns and anticipatory neural activity related to swallowing, which differ from those associated with chewing(Saito et al., 2024). The results underscore EEG’s value in disambiguating semantically similar actions when vision is limited.
Limitations of Integrating EEG: Despite these advantages, EEG does not always lead to improved recognition. Fig.5b presents failure cases of the multimodal model. The participant is actually operating PowerPoint, but the model incorrectly identifies the task as using a drawing application. Possible reason is that the participant is creating multiple rectangles, evoking visuomotor activity patterns that resemble those during freehand drawing. Prior studies have shown that such overlapping cognitive strategies lead to similar EEG signatures(Dvorak et al., 2018), making semantic discrimination harder.